OpenStudy (anonymous):

Use identities to find sinθ given secθ=5/2 and θ is in quadrant 4?

OpenStudy (anonymous):

\[\sec \theta = 1/ \cos \theta\] therefore \[\cos \theta = 2/5\]

OpenStudy (anonymous):

sin@= -2/5